OpenStudy (anonymous):

x - 0.1x + 0.06(x - 0.1x) = 70.79

OpenStudy (cwrw238):

x - 0.1x + 0.06(x - 0.1x) = 70.79 x - 0.1x + 0.06x - 0.006x = 70.79 1.06x - 0.106x = 70.79 0.954x = 70.79 now divide each side by 0.954 to get value of x